OpenStudy (anonymous):

i need a mathematical proof of x^0=1

myininaya (myininaya):

\[x^0=x^{n-n}=\frac{x^n}{x^n}=1\]

myininaya (myininaya):

as long as x does not =0

myininaya (myininaya):

see my use of law of exponents! :)

OpenStudy (anonymous):

is that an mathematical proof?

myininaya (myininaya):

Proof: Assume \[x \neq 0\] \[x^0=x^{n-n}\] since n-n=0 \[x^{n-n}=\frac{x^n}{x^n}\] by law of exponents \[\frac{x^n}{x^n}=x^nx^{-n}\] by law of exponents \[x^{n}x^{-n}=1\] since \[x^n, x^{-n}\] are multiplicative inverses af each other so we have thus proved \[x^0=1\].//

OpenStudy (anonymous):

what about 0/0?

myininaya (myininaya):

notice i said assume x does not equal 0

myininaya (myininaya):

0/0 is indeterminate form and so is 0^0

OpenStudy (anonymous):

thanks :)

OpenStudy (amistre64):

\[\lim_{x->0}x^0=1\]

OpenStudy (amistre64):

its like haveing online turrets :)

OpenStudy (amistre64):

tourette ....